# Monogamy of nonlocality from multipartite information causality

### v1

@@ Abstract
In this section, we present the proof of Lemma 1. Specifically, we demonstrate that the slice (9) is optimal for retrieving monogamy relations of the form (2) from bipartite and multipartite criteria (6),(7),(11) for information causality ($\mathcal{IC}$) in a tripartite $\mathcal{A},\mathcal{B},\mathcal{E}$ binary input $x,y,z\in\{0,1\}$ and binary output $a,b,e\in\{0,1\}$ Bell scenario.
Recall that, our aim is to retrieve upper bounds on the value CHSH expression $\beta(\mathcal{B},\mathcal{E})$ between $\mathcal{B}$$\mathcal{E}$ as a function of the (for a given) value of CHSH expression $\beta(\mathcal{A},\mathcal{B})$ between Alice and Bob. To this end, we observe that $\beta(\mathcal{A},\mathcal{B})$ and $\beta(\mathcal{B},\mathcal{E})$ are individually maximized by their respective PR-boxes $\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}},\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}$ where $L_{\mathcal{A}},L_{\mathcal{E}}$ are some local distributions for $\beta(\mathcal{A},\mathcal{E})$, respectively, such that,
$p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}}(a,b,e|x,y,z)=\frac{1}{2}\delta_{a\oplus b,xy}p_{L_{\mathcal{E}}}(e|z),$ (21)
$p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}}(a,b,e|x,y,z)=\left(\frac{1}{2}\delta_{b\oplus e,yz}\right)p_{L_{\mathcal{A}}}(a|x),$ (22)
@@
respectively. We note here that the product structure of these correlations follows from the respective no-signaling conditions. Specifically, the no-signaling condition forces the tripartite distribution to factorize whenever any two of the three parties share a PR-box.
As, in the tripartite binary input and binary output Bell scenario, any other extremal nonlocal tripartite no-signaling box cannot attain the maximal violation of the CHSH inequalities [34], its contribution can be effectively ignored. Consequently, without loss generality, we can restrict ourselves to considering a correlation $\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ formed a convex combination of correlations $\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}$ (21), $\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}$ (22) and a product distribution $\mathbb{P}_{\mathcal{A}}\otimes\mathbb{P}_{\mathcal{B}}\otimes\mathbb{P}_{\mathcal{E}}$ composed of the marginal distributions such that, $\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}=\alpha\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}+\gamma\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}+(1-\alpha-\gamma)\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}}\otimes\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{B}}\otimes\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{E}}$, where the coefficients $\alpha,\gamma\geq 0$, $\alpha+\gamma\leq 1$, such that,
$p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)=\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}}(a,b,e|x,y,z)+(1-\alpha-\gamma)p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}}}(a|x)p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{B}}}(b|y)p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{E}}}(e|z),$ (23)
where,
$p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}}}(a|x)=\frac{1}{\alpha+\gamma}\sum_{b,e\in\{0,1\}}\left(\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}}(a,b,e|x,y,z)\right)=\frac{1}{2},$ (24)
$p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{B}}}(b|y)=\frac{1}{\alpha+\gamma}\sum_{a,e\in\{0,1\}}\left(\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}}(a,b,e|x,y,z)\right),$ (25)
$p_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{E}}}(e|z)=\frac{1}{\alpha+\gamma}\sum_{b,e\in\{0,1\}}\left(\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}}(a,b,e|x,y,z)\right).$ (26)
Consequently, we have a family of two-parameter slices of tripartite no-signaling polytope characterized by the local distributions $L_{\mathcal{A}},L_{\mathcal{E}}$. In particular, in each of these slices, the value of the CHSH expressions are determined exclusively by the coefficients ${(\alpha,\gamma)}$ such that,
$\beta(\mathcal{A},\mathcal{B})=\frac{1+\alpha}{2};$ (27)
$\beta(\mathcal{B},\mathcal{E})=\frac{1+\gamma}{2}.$ (28)
Next, we demonstrate that for retrieving monogamy relations of the form (2) for bipartite (6),(7) and multipartite criteria (8),(11) for $\mathcal{IC}$, it is optimal to take the local distributions $L_{\mathcal{A}},L_{\mathcal{E}}$ to be white noise distributions $L_{\mathcal{A}}=\text{\text{WN}}_{\mathcal{A}},L_{\mathcal{E}}=\text{WN}_{\mathcal{E}}$, where,
$p_{\text{WN}_{\mathcal{A}}}(a|x)=\frac{1}{2}$ (29)
$p_{\text{WN}_{\mathcal{E}}}(e|z)=\frac{1}{2}.$ (30)
To prove the desired thesis, as a starting point, let us consider a tripartite distribution ${\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}$ which satisfies the bipartite $\mathcal{IC}$ criterion (7), such that,
@@
$\sum_{i=1}^{n}I_{{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}:G_{i},M^{\prime})+\sum_{i=2}^{n}I_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(X_{1}:X_{i}|G_{i},M^{\prime})\leq C+\sum_{i=2}^{n}H(X_{i})-H(X_{1},...,X_{n}),$ (31)
and/or multipartite (8) $\mathcal{IC}$ criterion,
$\sum_{k,i}I_{{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},\mathbf{M^{\prime}},G_{i})\leq\sum_{k}^{N-1}C_{k}+\sum_{i}^{n}I(X_{i+1}^{k},\dots,X_{n}^{k}:X_{i}^{k}),$ (32)
where we have used the respective protocols described in the main text such that mutual information terms $I_{{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}$ on the left hand side of (31),(32) depend only on the tripartite correlation $\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$. Notice that terms on right hand side of (31),(32) depend only on specifics of the communication task and are independent of the the tripartite correlation $\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ such that they are already equal to the terms on the right hand side of (7),(8). Let us now consider a distribution ¯$\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ obtained from the given distribution $\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ by flipping all outputs, such that,
$p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(a,b,e|x,y,z)=\alpha p_{\text{PR}_{\mathcal{A}\mathcal{B}}\otimes\bar{L}_{\mathcal{E}}}(a,b,e|x,y,z)+\gamma p_{\text{PR}_{\mathcal{B}\mathcal{E}}\otimes\bar{L}_{\mathcal{A}}}(a,b,e|x,y,z)+(1-\alpha-\gamma)p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}}}(a|x)p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{B}}}(b|y)p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{E}}}(e|z),$ (33)
where,
$p_{\bar{L}_{\mathcal{E}}}(e|z)=p_{{L}_{\mathcal{E}}}(e\oplus 1|z),$
$p_{\bar{L}_{\mathcal{A}}}(a|x)=p_{{L}_{\mathcal{A}}}(a\oplus 1|x),$
$p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{B}}}(b|y)=\frac{1}{2},$
$p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}}}(a|x)=p_{{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}}}(a\oplus 1|x),$
$p_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{E}}}(e|z)=p_{{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{E}}}(e\oplus 1|z).$ (34)
As flipping the outputs is a simple local post-processing, from the data processing inequality (12) we have the following inequality for terms on the left-hand side of (7),
@@
$\sum_{i=1}^{n}I_{{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}:G_{i},M^{\prime})+\sum_{i=2}^{n}I_{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(X_{1}:X_{i}|G_{i},M^{\prime})\leq\sum_{i=1}^{n}I_{{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}:G_{i},M^{\prime})+\sum_{i=2}^{n}I_{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}(X_{1}:X_{i}|G_{i},M^{\prime}),$ (35)
and similarly for the terms on the left-hand side of (8),
$\sum_{k,i}I_{{\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},\mathbf{M^{\prime}},G_{i})\leq\\ \sum_{k,i}I_{{\mathbb{P}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}}}(X_{i}^{k}:X_{i}^{1},\dots,X_{i}^{k-1},X_{i}^{k+1},\dots,X_{i}^{N-1},\mathbf{M^{\prime}},G_{i}),$ (36)
while terms on the right-hand side of (7) and (8) depend only on the channel capacity and the prior distribution of the inputs, hence, remain unaltered. Hence, given a distribution which satisfies (7) and/or (8) ${\mathbb{P}}_{\mathcal{A}\mathcal{B}\mathcal{E}}$, the flipped-outcomes distribution ¯$\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ also satisfies bipartite (7) and/or multipartite (8) formulations of $\mathcal{IC}$, respectively. Note, that the value of the CHSH expression, $\beta(\mathcal{A},\mathcal{B})$ and $\beta(\mathcal{B},\mathcal{E})$ remain unaltered when all outcomes are simultaneously flipped, as they only depend on the coefficients $(\alpha,\gamma)$ (27) (28).
Now, let us consider another tripartite distribution ~$\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ obtained by mixing equal proportions of the original tripartite distribution ${\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$ and the flipped-outcomes distribution ¯$\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}$, i.e., ~$\tilde{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}=\frac{1}{2}({\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}}+\bar{\mathbb{P}}^{(\alpha,\gamma)}_{\mathcal{A}\mathcal{B}\mathcal{E}})$, such that,

### v2

@@ Abstract
In this section, we prove Lemma 1. In particular, we show that the slice defined in Eq.(9) is optimal for deriving monogamy relations of the form (2), based on the bipartite and multipartite information causality ($\mathcal{IC}$) criteria given in Eqs. (6), (7), and (11), within the tripartite Bell scenario involving parties $\mathcal{A}$, $\mathcal{B}$, and $\mathcal{E}$, each with binary inputs $x,y,z\in{0,1}$ and binary outputs $a,b,e\in{0,1}$.
First of all, we start by recalling the depolarization procedure described in [40, 37], which states that any bipartite distribution $\mathbb{P}_{\mathcal{A},\mathcal{B}}$ can be transformed into the isotropic distribution using local operations and shared randomness (LOSR), keeping the CHSH value. That is,
[…]
$\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}}}$ $\quad\>=\quad\quad\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}}^{\text{iso}(\alpha)}}=\frac{1+\alpha}{2},$ (21b)
where $\alpha\in[0,1]$, $\text{PR}_{\mathcal{A}\mathcal{B}}$ is the PR-box, achieving the algebraic maximum value of the CHSH expression, $\beta(\mathcal{A},\mathcal{B})$ (i.e., $p_{\text{PR}_{\mathcal{A}\mathcal{B}}}(a,b|xy)=(1/2)\delta_{a\oplus b,xy}$), and $\text{W}_{\mathcal{A}\mathcal{B}}$ denotes the white noise distribution (i.e., $p_{\text{W}_{\mathcal{A}\mathcal{B}}}(a,b|xy)=1/4$).
[…]
$\sum_{\mathcal{E}}\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}}=\alpha\text{PR}_{\mathcal{A}\mathcal{B}}+(1-\alpha)\text{W}_{\mathcal{A}\mathcal{B}}.$ (22)
Here, $\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}$ refers to the CHSH value obtained by marginalizing the tripartite distribution over the party $\mathcal{E}$. Since the PR-box is an extremal point in the tripartite Bell scenario with binary inputs and outputs [31], the isotropic distribution $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}}$ must be a convex combination of the PR-box and another distribution $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime}$ satisfying the condition (22). From (21), we should then have:
[…]
$\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}}$ $\quad\>=\quad\quad\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}(\alpha)}}=\frac{1+\alpha}{2},$ (23b)
@@
where $L_{\mathcal{E}}$ denotes some local distribution for the party $\mathcal{E}$, and $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime}$ is any tripartite distribution addressing (22), i.e., $\sum_{\mathcal{E}}\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime}=W_{\mathcal{A}\mathcal{B}}$, such that $\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime}}=1/2$. Notice that $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime}$ might, in principle, be a non-local correlation. Importantly, the product structure of the correlation $\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}$ follows from the respective no-signaling conditions: Specifically, the no-signaling condition forces the tripartite distribution to factorize whenever any two of the three parties share a PR-box [24].
[…]
@@
$\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}(\alpha)}}$ $\quad\>=\quad\quad\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}},$ (24b)
@@
$\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}(\alpha)}}$ $\quad\>=\quad\quad\beta(\mathcal{A},\mathcal{B})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}},$ (24c)
where $\gamma\in[0,1]$, $L_{\mathcal{A}}^{\prime}$ is a local distribution for $\mathcal{A}$, and $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime(\alpha)}$ is any tripartite distribution such that $\sum_{\mathcal{A}}\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime(\alpha)}=W_{\mathcal{B}\mathcal{E}}$, and $\beta(\mathcal{B},\mathcal{E})_{\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}}=1/2$. According to (24c), the transformation $\text{LOSR}_{\mathcal{B}\mathcal{E}}$ does not affect the contribution $\alpha\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}$ from (23). At most, this component might be relabeled into another version of the PR-box, still achieving the maximal value for some relabeled CHSH inequality. Hence, we may write $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime(\alpha)}$ in (24a) in terms of one more parameter $\epsilon\in[0,1]$, such that:
$\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime(\alpha,\epsilon)}=\epsilon\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}+(1-\epsilon)\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}.$ (25)
$\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}$ should address the constraints of $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime(\alpha,\epsilon)}$ and $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime}$ in (23) and (24), respectively. i.e., with (25) in (24a) we have,
$\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\text{iso}^{\prime}(\alpha,\gamma)}=\alpha\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}+\gamma\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}^{\prime}+(1-\alpha-\gamma)\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}.$ (26)
and $\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}$ should respect,
$\sum_{\mathcal{E}}$ $\left\{\gamma\text{PR}_{\mathcal{B}\mathcal{E}}\otimes L_{\mathcal{A}}^{\prime}+(1-\gamma)\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}\right\}=W_{\mathcal{A}\mathcal{B}},$ (27a)
$\sum_{\mathcal{A}}$ $\left\{\epsilon\text{PR}_{\mathcal{A}\mathcal{B}}\otimes L_{\mathcal{E}}+(1-\epsilon)\mathbb{P}_{\mathcal{A}\mathcal{B}\mathcal{E}}^{\prime\prime\prime}\right\}=W_{\mathcal{B}\mathcal{E}}.$ (27b)
Notice that in (26) we considered the fact $\alpha=(1-\gamma)\epsilon$. At this point, it follows as a corollary from the depolarization procedure (21) that tripartite correlations of binary inputs and binary outputs Bell scenarios can be transformed to the distribution (26), without changing the CHSH values, $\beta(\mathcal{A},\mathcal{B})$ and $\beta(\mathcal{B},\mathcal{E})$. In particular, in each of these slices, the value of the CHSH expressions are determined exclusively by the coefficients $(\alpha,\gamma)$} such that,